in April 2017,Reinko purchased and placed a rental house in service.She paid 184000, including 27000 for the land.Compute her 2017 depreciation on the rental.
Answer #1

Ans- Cost basis of rental property for depreciation = 184,000-27,000= 157,000

Depreciation percentage for 2017 using straight line depreciation with a 27.5 year classification= 2.576%

Depreciation expense= 157,000*2.576%= 4,044